OpenStudy (anonymous):

prove 1-sinᶿ/cosᶿ=cosᶿ/1-sinᶿ...

OpenStudy (anonymous):

is this \[1-\frac{\sin(\theta)}{\cos(\theta)}\]

OpenStudy (anonymous):

good question, it's not true if you take the question to be \[(1- \sin)/(\cos) = \cos/(1-\sin)\]

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

\[1-\frac{\sin(\theta)}{\cos(\theta)}=\frac{\cos(\theta)-\sin(\theta)}{\cos(\theta)}\] \[=\frac{\cos(\theta)-\sin(\theta)}{\cos(\theta)}\times \frac{\cos(\theta)}{\cos(\theta)}\] \[=\frac{\cos^2(\theta)-\cos(\theta)\sin(\theta)}{cos^2(\theta)}\] \[\frac{\cos(\theta)(1-\sin(\theta)}{1-sin^2(\theta)}\] \[=\frac{\cos(\theta)(1-\sin(\theta)}{(1+\sin(\theta))(1-\sin(\theta))}\] \[=\frac{\cos(\theta)}{1+\sin(\theta)}\]

OpenStudy (anonymous):

which is not exactly what you posted but i think it is right.